OpenStudy (jjuden):

Determine the prime factorization of 90

OpenStudy (jjuden):

a 2 • 3^2• 5 b 2^2• 3 • 5 c 2^3• 11 d 2^2• 11

OpenStudy (shad0w):

Break it down to smaller numbers: 90 = 9 * 10, so prime-factor 9 and 10. 90 = 9 * 10 = 3 * 3 * 2 * 5 = 2 * 3^2 * 5 So the prime factors of 90 are 2, 3 (twice, or squared), and 5.